OpenStudy (anonymous):

Point X is located at (2, −6) and point Z is located at (0, 5). Find the y value for the point Y that is located one over five the distance from point X to point Z. A) −3.6 B) −3.8 C) −4 D) −4.2

OpenStudy (sleepyhead314):

mmm I think you would find the distance between the two given points first... \[distance formula: \sqrt{(x _{2}-x _{1})^{2} + (y _{2}-y _{1})^{2}}\]when given points (x1, y1) and (x2, y2)

OpenStudy (anonymous):

I tried that but I couldn't find the right one

OpenStudy (sleepyhead314):

hmmm, what'd you get when you tried that? :)

OpenStudy (anonymous):

2.7 idk what I did wrong

OpenStudy (sleepyhead314):

aww I'm getting stuck too :/ @ganeshie8 @Luigi0210 @ParthKohli could you help if you're not too busy? :3

OpenStudy (sleepyhead314):

mmmm the distance between the two y values is 5 - (-6) = 11 right? and one over five of 11 = 2.2 so from -6 + 2.2 = ? <-- that might be it?

OpenStudy (sleepyhead314):

@lucy27 you get it?

OpenStudy (anonymous):

you have points (2,-6) and (0,5) and you're trying to find what is a fifth of that distance. \[\sqrt{(x_{2}-x _{1})+(y _{2}-y _{1})}\] \[\sqrt{(-2)+(5+6)}\] \[\sqrt{-2+11}\] \[\sqrt{9}\] =3. now what's 1/5 of 3? 1/5x3 =3/5 =0.6. this is what i got?

OpenStudy (anonymous):

oh oops i did the wrong thing!! here i'll do the right one..

OpenStudy (anonymous):

you know the midpoint formula right? M=(x1+x2)/(y1+y2)/2. but what if, instead of finding the centre (or the half, which is /2) we would find the FIFTH, so intead of /2 we would do /5?

OpenStudy (anonymous):



OpenStudy (anonymous):

Midpt: \[\frac{ 2+0 }{ 5 }, \frac{ -6+5 }{ 5 }\] 2/5, -1/5 0.4, -0.2<--your answer i think.prob wrong though...this is a hard question..

OpenStudy (sleepyhead314):

the distance between the two y values is 5 - (-6) = 11 right? and one over five of 11 = 2.2 so from -6 + 2.2 = ? <-- that might be it? I got -3.8 B